needing help...
divide 40x/64x by 5x/8y

Answer by checkley71(8403)

(40X/64X)/5X/8Y OR 40X*8Y/64X*5X OR 320XY/320X~2 OR Y/X
I'LL TRY ANOTHER APPROACH
40X/64X=40/64=5/8 NOW TO DIVIDE THIS FRACTION BY ANOTHER FRACTION THE SECOND FRACTION MUST BE INVERTED OR 8Y/5X NOW YOU NEED TO MULIPLY THE TWO FRACTION
(5/8)/(8Y/5X) THE 8s CANCEL & THE 5s CANCEL LEAVING THE FRACTION Y/X